OpenStudy (anonymous):

I WILL GIVE A MEDAL!!! Divide: (x+2/2x)/(6/x+3) Discuss why the degree of the resulting denominator did not change from your expression’s degree.

OpenStudy (anonymous):

\[\huge \frac{ x+\frac{ 2 }{ 2x } }{ \frac{ 6 }{ x+3 } }\] is this your problem?

OpenStudy (freckles):

choice A \[\frac{x+\frac{2}{2x} }{\frac{6}{x}+3}\] choice B \[\frac{\frac{x+2}{2x} }{\frac{6}{x}+3}\] choice C \[\frac{\frac{x+2}{2x} }{\frac{6}{x+3}}\]

OpenStudy (anonymous):

choice C

OpenStudy (anonymous):

\[\huge \frac{\frac{x+2}{2x} }{\frac{6}{x+3}}\] ok cool, so have you tried anything yet?

OpenStudy (anonymous):

Nope I am lost

OpenStudy (jhannybean):

Which part are you confused with?

OpenStudy (anonymous):

The whole thing. I'm not that good with dividing these things.

OpenStudy (anonymous):

\[\huge \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } \implies \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

okay so I would put it as this: